OpenStudy (katherinesmith):

What is the range of y = tan(x) + 2?

OpenStudy (katherinesmith):

my guess is all real numbers > -2. is this correct?

OpenStudy (katherinesmith):

@phi

OpenStudy (katherinesmith):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

the range of y = tan(x) is the set of all real numbers this is also true for y = tan(x) + 2 because shifting it up or down doesn't change the range

jimthompson5910 (jim_thompson5910):

you'll be able to see it with a graph

OpenStudy (katherinesmith):

i thought the domain was all real numbers but the range was different

jimthompson5910 (jim_thompson5910):

the domain is not the set of real numbers, there are restrictions

jimthompson5910 (jim_thompson5910):

there are values of x you cannot use

OpenStudy (katherinesmith):

so what would the domain be

OpenStudy (phi):

doesn't tan go from -inf to + inf ? adding 2 does not change that

jimthompson5910 (jim_thompson5910):

hint: tan(x) = sin(x)/cos(x)

OpenStudy (katherinesmith):

okay, i see what you mean. so how do i figure the restrictions of the domain to find the domain of the function

jimthompson5910 (jim_thompson5910):

use my hint and remember that you cannot divide by zero

OpenStudy (katherinesmith):

is 0 the only restriction?

jimthompson5910 (jim_thompson5910):

no x = 0 is allowed

OpenStudy (katherinesmith):

oh my god

jimthompson5910 (jim_thompson5910):

tan(0) = 0, so that's defined

OpenStudy (phi):

see a plot https://www.wolframalpha.com/input/?i=plot++tan+x

OpenStudy (katherinesmith):

then how the heck do i find the domain.

OpenStudy (anonymous):

\[ \cos(x)\neq 0 \]

jimthompson5910 (jim_thompson5910):

you need to find out when cos(x) = 0 is true

jimthompson5910 (jim_thompson5910):

that way, you'll know which values of x to exclude from the domain

OpenStudy (katherinesmith):

how do i find that out...

jimthompson5910 (jim_thompson5910):

use the unit circle

jimthompson5910 (jim_thompson5910):

look for points on the unit circle that have an x coordinate of 0 then record the angle

OpenStudy (katherinesmith):

90 degrees and 270 degrees

jimthompson5910 (jim_thompson5910):

so this means that x cannot be 90 or x cannot be 270 in general, x cannot be equal to 90+180n where n is any integer

jimthompson5910 (jim_thompson5910):

so the domain is the set of all real numbers x, but x cannot equal 90+180n where n is any integer

jimthompson5910 (jim_thompson5910):

this is assuming x is in degrees

OpenStudy (katherinesmith):

okay. thank you for the help

jimthompson5910 (jim_thompson5910):

yw